Use POLYA'S FOUR-STEP PROBLEM SOLVING STRATEGY to solve the problems.

  1.  Layla is going to buy 30 multi-vitamin capsules, some P10 and some P18. If she has P340, what is the maximum number of P18 multi-vitamin capsules she can buy? (HINT: Use inequality)
  2. Yesterday, Ling got confused about which day of the week it was. Whenever they go to long vacation, he forgets. Johnson said its Friday, but Roger said its Saturday. Ling asked what day is it tomorrow? Roger said its Monday, while Johnson said its Tuesday. Ling asked again, what day has it yesterday? Johnson said Wednesday, while Roger said its Thursday. Johnson and Roger have given one correct answer and two wrong answers. What day is it today? 
1
Expert's answer
2021-05-07T09:29:42-0400

1.

x - P10 capsules, y - P18 capsules

"x+y\le30"

"10x+18y\le340\implies 5x+9y\le170"

Objective: "max \ y"


We have three points for considering:

"(30,0),\ (0,170/9),\"and:

"x=30-y\implies 5\cdot(30-y)+9y=170"

"4y=20\implies y=5,\ x=25" ,

and the third point: "(25,5)"

"max\ y=18.9"

Answer: "18"


2.

One correct answer: Thursday (yesterday) is before Friday (today).

Answer: Friday